OpenStudy (raffle_snaffle):

You have 357 mL of chlorine trifluoride gas at 699 mmHg and 45 * C. What is the mass in grams of the sample?

OpenStudy (raffle_snaffle):

@ganeshie8

OpenStudy (raffle_snaffle):

what the heck is mmHg?

OpenStudy (raffle_snaffle):

why is that a pressure?

OpenStudy (aric200):

mmHg is A millimeter of mercury is a manometric unit of pressure, formerly defined as the extra pressure generated by a column of mercury one millimetre high. It is now defined precisely as 13.5951 × 9.80665 = 133.322387415 pascals. Wikipedia

OpenStudy (raffle_snaffle):

okay thank you

OpenStudy (aric200):

MM is millimeter and Mercury on the Periodic table is Hg

OpenStudy (raffle_snaffle):

so can you plz point me i the right direction as to how to solve the problem?

OpenStudy (aric200):

Hold on a second because I have not even read it yet :P

OpenStudy (raffle_snaffle):

wait we use pv=nRT

OpenStudy (raffle_snaffle):

all i need to do is convert mL to L and C to K then I can apply pv=nRT

OpenStudy (aric200):

P = 699mm Hg = 0.920atm 699/760atm = V = 427mL = .427 L = 45 T = 45oC +273 k = 318K R = .0821Latm/mol.K find n number of moles n = PV / RT = .920 X .427 / (.0821 X318) = 0.0150 mass = molesX molarmass = 0.0150 X 92.45 = 1.39g

OpenStudy (raffle_snaffle):

once i obtain the mols of the compound I can solve for grams of the compound

OpenStudy (aric200):

I honestly dont know how to solve this all I did was google search it because I was lost

OpenStudy (aric200):

Sorry :/

OpenStudy (raffle_snaffle):

what? lost?

OpenStudy (aric200):

Yeah Im not really sure how to solve this?

OpenStudy (abmon98):

PV=nRT Pressure is measured in pascal. you need to know that 1mmhg is equivalent to 133.322368pascal, if we convert 699mmhg to pascal.we multiply 699 by 133.322368 to get 93192.335232 pascals. The SI unit for volume is m^3. 1ml is 1*10^-6. so changing the 375 ml to m^3 we multiply by 10^-6 getting 375*10^-6 m^3. R is a gas constant value of 8.31 and temperature has to be measure in kelvin. to change from degrees we add 273 to 45. Number of moles=mass/molar mass

OpenStudy (abmon98):

93192.335232*375*10^-6=mass/35.5*8.31*318 35.5(93192.335232*375*10^-6)=mass*2642.58 mass=1240.622/2642.58